• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Iodice

    Log entry time 20:49:36 on January28,2004

    Entry number 116980

    keyword=Q3 VACUUM LEAK

    The left Q3 magnet has a vacuum leak problem that could not be fixed
    today by normal turbo pump down procedures. Ed et al. were working
    the all day to try to fix it.
    They will continue to work tomorrow morning. At the moment there is no estimate for when Q3 will be recovered.

    The present SWING shift and next Jan. 29 OWL shift HAVE BEEN CANCELLED. Jan. 29 DAY shift is in STANDBY.